# A Measure for Transparent Comparison of Linguistic Diversity in Multilingual NLP Data Sets

### v1

@@ 7 Conclusion
We have shown that the linguistic diversity of NLP data sets can be consistently assessed by two independent measures, TI (proposed in previous work) and J (proposed in this paper). Both of these measures show that a high number of languages and language families included in a data set is not sufficient to ensure linguistic diversity.
@@
An advantage of the J score over TI and other previous indicators of linguistic diversity is its capacity to show what kinds of languages are missing in a given data set in comparison to a reference. Assessing popular NLP data sets with this measure revealed that the most underrepresented languages are those with rich morphology. This kind of direct and transparent comparison can improve multilingual NLP coverage in the long run.
@@
Both measures of morphological diversity that we propose rely on text features (the mean word length). Although we show that the mean word length is strongly correlated with an independent measure of morphological complexity (WALS). it remains an aggregated measure (one feature per language). For an even more transparent measure of linguistic diversity, it would be desirable to obtain more nuanced morphological features.
Another limitation of our study is that we do not propose syntactic features that could be extracted from text. We focused here on the current gap in the available linguistic features (the lack of morphological features in lang2vec), but devising text-based syntactic features would deserve more attention in future work.
@@ Appendix B Word length and morphological complexity
| ISO396-3 | MWL | |
| abk | 7.17 | 0.62 |
| apu | 7.67 | 0.60 |
| arz | 4.44 | 0.49 |
| bsn | 6.02 | 0.69 |
| ckt | 8.45 | 0.50 |
| deu | 4.87 | 0.55 |
| ell | 4.72 | 0.53 |
| eng | 4.18 | 0.42 |
| eus | 5.70 | 0.64 |
| fin | 6.23 | 0.66 |
| fra | 4.41 | 0.45 |
| hae | 5.91 | 0.53 |
| hau | 4.08 | 0.38 |
| heb | 3.94 | 0.54 |
| ind | 5.42 | 0.40 |
| kan | 5.22 | 0.65 |
| kat | 4.78 | 0.50 |
| khk | 5.66 | 0.53 |
| kut | 4.60 | 0.37 |
| lvk | 4.77 | 0.67 |
| qvi | 8.18 | 0.71 |
| rus | 4.79 | 0.52 |
| spa | 4.37 | 0.45 |
| swh | 5.72 | 0.71 |
| tur | 6.07 | 0.76 |
| vie | 3.20 | 0.21 |
| yaq | 5.31 | 0.57 |
| yor | 3.52 | 0.25 |
@@ Appendix C Word Length Adjustments for Logographic Scripts
Name and main references |
N(L) | N(F) | TI | TI | ||
Universal Dependencies (UD) |
106* | 20* | 0.567 | 0.736 | 0.337 | 0.665 |
Bible 100 |
103* | 30* | 0.649 | 0.811 | 0.302 | 0.617 |
mBERT |
97* | 15* | 0.559 | 0.710 | 0.316 | 0.617 |
XTREME |
40 | 14 | 0.612 | 0.775 | 0.311 | 0.471 |
XGLUE |
19 | 7* | 0.517 | 0.674 | 0.297 | 0.580 |
XNLI |
15 | 7* | 0.557 | 0.711 | 0.321 | 0.704 |
XCOPA |
11 | 11 | 0.586 | 0.737 | 0.336 | 0.634 |
TyDiQA |
11 | 10 | 0.626 | 0.751 | 0.343 | 0.552 |
XQuAD |
12* | 6* | 0.523 | 0.680 | 0.318 | 0.634 |
TeDDi |
89 | 51 | 0.706 | - | 0.361 | - |
In the case of logographic scripts, we scale the observed word length proportionally to the difference between Chinese original script and Pinyin so that the scaled length is comparable to alphabetic scripts. We estimate the scalar as a function of the character-level type-token ratio on a sample of text:
Where is a constant representing the mean word length ratio between Pinyin and the original Chinese script, is the language in question and is the character-level type-token ratio on a text sample of the length of 500 word-level tokens in a given language. The scalar should be applied only if its value is greater than 1.
Table 5 shows revised diversity scores after the mean word length adjustments for logographic scripts. In our study, only three languages Chinese, Japanese and Korean required adjustments.

### v2

@@ 8 Conclusion
We have shown that the linguistic diversity of NLP data sets can be consistently assessed by two independent measures, TI (proposed in previous work) and Jmm (proposed in this paper). Both of these measures show that a high number of languages and language families included in a data set is not sufficient to ensure linguistic diversity.
@@
An advantage of the Jmm score over TI and other previous indicators of linguistic diversity is its capacity to show what kinds of languages are missing in a given data set in comparison to a reference. Assessing popular NLP data sets with this measure revealed that the most underrepresented languages are those with rich morphology. This kind of direct and transparent comparison can improve multilingual NLP coverage in the long run.
Acknowledgements
This research is partially supported by the Swiss National Science Foundation (SNSF) grants 176305 and PCEFP1_186841. We thank the anonymous reviewers for their suggestions, which have improved the clarity of the paper.
@@ Acknowledgements
A limitation of our study is that the two levels of linguistic structures are represented with different features: syntax with expert features from linguistic data bases and morphology with mean word length as a text feature. Our results suggest that the two measures agree more at the level of syntax than at the level of morphology. To draw sound conclusions about the impact of the structural level on the agreement between the two measures, we would need both kinds of features for both levels. While we indirectly compare text and expert features at the level of morphology (via the correlation test), we do not propose syntactic features that could be extracted from text. We focused here on the current gap in the available linguistic features (the lack of morphological features in lang2vec), but devising and validating text-based syntactic features would deserve more attention in future work.
@@ Appendix B Word length and morphological complexity
ISO396-3 | Name | MWL | |
abk | Abkhazian | 7.17 | 0.62 |
apu | Apurinã | 7.67 | 0.60 |
arz | Egyptian Arabic | 4.44 | 0.49 |
bsn | Barasana-Eduria | 6.02 | 0.69 |
ckt | Chukchi | 8.45 | 0.50 |
deu | German | 4.87 | 0.55 |
ell | Modern Greek | 4.72 | 0.53 |
eng | English | 4.18 | 0.42 |
eus | Basque | 5.70 | 0.64 |
fin | Finnish | 6.23 | 0.66 |
fra | French | 4.41 | 0.45 |
hae | Eastern Oromo | 5.91 | 0.53 |
hau | Hausa | 4.08 | 0.38 |
heb | Modern Hebrew | 3.94 | 0.54 |
ind | Indonesian | 5.42 | 0.40 |
kan | Kannada | 5.22 | 0.65 |
kat | Georgian | 4.78 | 0.50 |
khk | Halh Mongolian | 5.66 | 0.53 |
kut | Kutenai | 4.60 | 0.37 |
lvk | Lavukaleve | 4.77 | 0.67 |
qvi | Imbabura Highland Quichua | 8.18 | 0.71 |
rus | Russian | 4.79 | 0.52 |
spa | Spanish | 4.37 | 0.45 |
swh | Swahili | 5.72 | 0.71 |
tur | Turkish | 6.07 | 0.76 |
vie | Vietnamese | 3.20 | 0.21 |
yaq | Yaqui | 5.31 | 0.57 |
yor | Yoruba | 3.52 | 0.25 |
@@ Appendix C Word Length Adjustments for Logographic Scripts
Name and main references | N(L) | N(F) | TIsyn | TImorph | ||
Universal Dependencies (UD) | 106* | 20* | 0.567 | 0.736 | 0.337 | 0.665 |
Bible 100 | 103* | 30* | 0.649 | 0.811 | 0.302 | 0.617 |
mBERT | 97* | 15* | 0.559 | 0.710 | 0.316 | 0.617 |
XTREME | 40 | 14 | 0.612 | 0.775 | 0.311 | 0.471 |
XGLUE | 19 | 7* | 0.517 | 0.674 | 0.297 | 0.580 |
XNLI | 15 | 7* | 0.557 | 0.711 | 0.321 | 0.704 |
XCOPA | 11 | 11 | 0.586 | 0.737 | 0.336 | 0.634 |
TyDiQA | 11 | 10 | 0.626 | 0.751 | 0.343 | 0.552 |
XQuAD | 12* | 6* | 0.523 | 0.680 | 0.318 | 0.634 |
TeDDi | 89 | 51 | 0.706 | - | 0.361 | - |